If the demand function for a good is Q=140 – 5P, what is the price elasticity of demand at P = 

15 rupees?



1
Expert's answer
2021-04-12T07:09:11-0400

"Q=140-5P, \space P=15 rupees"

"Q=140-5\times15=65\space rupees"

"E_d=\frac{dQ}{dP}\times \frac{P}{Q}=-5\times\frac{15}{65}=-1.15"